OpenStudy (jackie4uxd):

Find the absolute extrema of the function on the closed interval. Use a graphing utility to verify your results. h(t) = (t)/((t − 3)^2) , [4, 7] Absolute Max (?,?) Absolute Min (?,?)

myininaya (myininaya):

First step is to differentiate h.

OpenStudy (jackie4uxd):

using the quotient rule? which would be \[\frac{ -3 }{ (t-3) ^2}\]

myininaya (myininaya):

hmm.... this is the quotient rule applied to your h function: \[h'(t)=\frac{(t)'(t-3)^2-t[(t-3)^2]'}{[(t-3)^2]^2}\]

myininaya (myininaya):

I want to save you pain. Do not multiply anything out at this moment. Just differentiate those factors with the ' on them.

OpenStudy (jackie4uxd):

\[-\frac{t+3 }{(t-3)^3 }\]

myininaya (myininaya):

(t)'=1 [(t-3)^2]'=2(t-3)

myininaya (myininaya):

on you canceled a factor let me recheck it

myininaya (myininaya):

yeah still missing something

myininaya (myininaya):

no you did fine lol

myininaya (myininaya):

I wasn't expecting you to simplify but okay lol

myininaya (myininaya):

Ok and we have critical numbers t=-3 which doesn't happen in the interval [4,7]

myininaya (myininaya):

so the only thing you need to look at are endpoints

OpenStudy (jackie4uxd):

loso i have to plug in the endpoints and that will give me the values ?

myininaya (myininaya):

\[h(4)=? \\ h(7)=? \]

myininaya (myininaya):

yep

myininaya (myininaya):

If we had any critical numbers to occur in between 4 and 7 we would plug in those as well but we didn't

OpenStudy (jackie4uxd):

oh alrighty thank you !

myininaya (myininaya):

You got it from here?

OpenStudy (jackie4uxd):

i think so for h(4) i got -7 for h(7) i got um kinda weird -10/64?

myininaya (myininaya):

\[h(4)=\frac{4}{(4-3)^2} \\ h(7)=\frac{7}{(7-3)^2}\]

OpenStudy (jackie4uxd):

oh i plugged it in the derivative thats why

myininaya (myininaya):

The derivative was just to give us any critical numbers between 4 and 7 The function that is actually going to tell where (4,h(4)) and (7,h(7)) are going to be is h

OpenStudy (jackie4uxd):

i got for h(4)=4 and for h(7)=7/16

OpenStudy (jackie4uxd):

i got the absolute maximum right (4,4) but i got the absolute minimum wrong (7, 7/16)

myininaya (myininaya):

sounds good to me